OpenStudy (anonymous):

Find the arc length of x=1/3(y^2+2)^3/2 y=0 y=1

OpenStudy (anonymous):

apply formula and done :DDD

OpenStudy (anonymous):

I mean yeah I know that I'm just having problems simplifying it

OpenStudy (anonymous):

y=((3x)^2/3-2)^0.5

OpenStudy (anonymous):

How?

myininaya (myininaya):

did you find the derivative of your x with respect to y?

OpenStudy (anonymous):

first multiply it with 3 for getting the 1/3 from the left side to the right side. Second. take the 2/3 on both side so that the 3/2 power will be gone from the right side. Third, -2 on both side (what you are left with is (3x)^2/3-2=y^2) and Last sqrtroot on both sides.

myininaya (myininaya):

you don't have to solve for y

OpenStudy (anonymous):

y' = (2/3) (y^2+2)^(1/2) (2y)

OpenStudy (anonymous):

urg i meant *x' = *

OpenStudy (anonymous):

which simplifes to x' = (y/3) (y^2 + 2)^(1/2)

myininaya (myininaya):

\[\frac{dx}{dy}=\frac{1}{3} \frac{3}{2} (y^2+2)^\frac{1}{2} (2y)\]

OpenStudy (anonymous):

O.O forgot about the 3/2

OpenStudy (anonymous):

Lol yeah I noticed it should be 1/2(y^2+2)^1/2 (2y) right?

OpenStudy (anonymous):

essentially yes

OpenStudy (anonymous):

which can be reduced to y (y^2 + 2)^(1/2)

OpenStudy (anonymous):

then use L = ∫ sqrt(1 + (dx/dy)^2]. I believe you'll have a perfect square under the radical after you simplify.

OpenStudy (anonymous):

Okay let me try it out to see what I get :)

OpenStudy (anonymous):

Okay so I got up to sqrt 1+(y^2(y^2+2) dy, should I distribute the y^2?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

Okay and after that what do I do?

OpenStudy (anonymous):

did you get y^4 + 2y^2 + 1 ?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

ok, it's a perfect square

OpenStudy (anonymous):

So y^2+2y?

OpenStudy (anonymous):

no it's (y^2 + 1)^2

OpenStudy (anonymous):

since that's under the square root, the exponents cancel leaving you with ∫ (y^2+1)dy, from 0 to 1. This is just a simple integral.

OpenStudy (anonymous):

Is the integral 1/2[1/3u^3]?

OpenStudy (anonymous):

no, it's (1/3)y^3 + y

OpenStudy (anonymous):

evaluate the limit and you should get (1/3) + 1 = 4/3

OpenStudy (anonymous):

Oh lol I thought there was a squared there that's why :p

OpenStudy (anonymous):

No wait yeah there is a squared isn't there?

OpenStudy (anonymous):

it cancels with root in the formula

OpenStudy (anonymous):

Okay got it thank you so so much ^-^